OpenStudy (anonymous):

PLEASE TEACH ME HOW TO DO IT! A shirt is on sale for 30% off the regular price. The regular price is $27.00. What is the sale price? $8.10 $18.20 $18.80 $18.90

hartnn (hartnn):

30% of 27 = 30/100 * 27 = ... ?

OpenStudy (anonymous):

8.1 = 8.10

hartnn (hartnn):

yep ;)

OpenStudy (anonymous):

THANKS :D

OpenStudy (anonymous):

That is not the answer.

OpenStudy (anonymous):

???

hartnn (hartnn):

sorry!

hartnn (hartnn):

thats the amount you subtract! :P

hartnn (hartnn):

sale price = regular price - discount = 27- 8.1

OpenStudy (anonymous):

ook!

hartnn (hartnn):

thanks for pointing it out @EulersEquation

OpenStudy (anonymous):

Thanks both!

OpenStudy (anonymous):

Ok, this is pretty simple if you think about it. If you take 10% of 27.00 you get 2.7, right? So multiply that by 3 because 10 *3 = 30, and you get 8.1. Then subtract 27- 8.1 and you get your answer 18.9

OpenStudy (anonymous):

I like to work problems of this type as follows: 30% off means that I still have 70% left. 30% + 70% = 100%. So I multiply the original price by the percent that I have left: $27.00 * 70% then directly gives me the answer in one step instead of two!